OpenStudy (anonymous):

The measure of an angle is 3 more than twice its supplement. Find the measure of an angle .

OpenStudy (anonymous):

S1= 121 Check answer is 59-121= 180.

OpenStudy (anonymous):

Good job !

OpenStudy (anonymous):

@ferlauren thanks.

OpenStudy (pawanyadav):

Sum of two supplementry angle is 180° If you take one angle X then other will be 180°-X.

OpenStudy (pawanyadav):

Now apply the condition that is given in the question

OpenStudy (pawanyadav):

Can you do this?

OpenStudy (vats-sagar):

x=(2(180-x)+3) 3x=363 x=121

OpenStudy (pawanyadav):

I think we should have leave the answer's calculation for the asker